OpenStudy (anonymous):

find d/dx (ln (x-1/x^2+4))

OpenStudy (anonymous):

Split it up using Ln(a/b) = Ln(a) - Ln(b) and do as normal

OpenStudy (anonymous):

i dont know how to do any of that at all

OpenStudy (anonymous):

Then you are truly lost

OpenStudy (anonymous):

a=(x-1) b=x^2+4 am I right newton?